OpenStudy (becca9898):

I need help with factoring polynomials. Can someone help me with this problem? Determine two different values of “b” in x^2 + bx + 30 so that the expression can be factored into the product of two binomials. Explain how you determined those values and show each factorization.

OpenStudy (anonymous):

there are several possible answers. think of two numbers whose product is 30, then add them together

OpenStudy (anonymous):

for example \(15\times 2=30\) so you could choose \(b=15+2=17\) or you could use \(5\times 6=30\) so \(b=5+6=11\) there are more